Palencia Paints Corporation has a target capital structure of 35% debt and 65% common equity, with no preferred stock. Its before-tax cost of debt is 9%, and its marginal tax rate is 40%. The current stock price is P0 = $26.00. The last dividend was D0 = $2.75, and it is expected to grow at a 6% constant rate. What is its cost of common equity and its WACC? Round your answers to two decimal places. Do not round your intermediate calculations.
Answer #1

cost of common equity=(D1/Current price)+Growth rate

=[(2.75*1.06)/26]+0.06

=17.21%(Approx).

Cost of debt after-tax=9*(1-tax rate)

=9*(1-0.4)=5.4%

WACC=Respective costs*Respective weight

=(17.21*0.65)+(0.35*5.4)

=13.08%(Approx).